# Optimal Computation in Anonymous Dynamic Networks

### v7

@@ 1.1 Background
Dynamic networks. The study of theoretical and practical aspects of highly dynamic distributed systems has garnered significant attention in recent years [11, 40, 43]. These systems involve a constantly changing network of computational devices called processes (sometimes referred to as “processors” or “agents”). Pairs of processes may send each other messages only through links that appear or disappear unpredictably. This dynamicity is typical of modern real-world systems and is the result of technological innovations, such as the spread of mobile devices, software-defined networks, wirelesses sensor networks, wearable devices, smartphones, etc.
@@
Connected networks. There are several models of dynamicity [11]; a popular choice is the 1-interval-connected network model [38, 47]. Here, a fixed set of $n$ processes communicate through links forming a time-varying graph, i.e., a graph whose edge set changes at discrete time units called rounds (thus, the system is synchronous); such a graph changes unpredictably, but is assumed to be connected at all times.
@@
Disconnected networks. The 1-interval-connected network model may not be a suitable choice for many real systems, due to the very nature of dynamic entities (think of P2P networks of smart devices moving unpredictably) or due to transient communication failures, which may compromise the network’s connectivity. A more realistic assumption is that the union of all the network’s links across any $\tau$ consecutive rounds induces a connected graph on the processes [36, 44, 49]. [^1] It is well known that non-trivial (terminating) computation requires some conditions on temporal connectivity to be met, as well as some a-priori knowledge by processes (refer to Proposition 2.3). We say that such a network is $\tau$-union-connected, and the smallest such parameter $\tau\geq 1$ is called dynamic disconnectivity. [^2] Another widely used parameter for dynamic networks is the dynamic diameter $d$. The relationship between $\tau$ and $d$ is discussed at the end of Section 2.3. In particular, it is worth noting that every occurrence of the parameter $\tau$ in the running times of our algorithms can be safely replaced with $d$. Observe that 1-interval-connected networks can equivalently be characterized as 1-union-connected networks.
@@
Networks with unique IDs. A large number of research papers have considered dynamic systems where each process has a distinct identity (unique IDs[38]. In this setting, there are efficient algorithms for consensus [39], broadcast [10], counting [38, 47], and many other fundamental problems [37, 43].
@@
It should be noted that networks with unique IDs allow for very simple algorithms for a wide variety of problems. Indeed, in a 1-interval-connected network of $n$ processes, if each process broadcasts the set of its “known IDs” at every round (initially, a process only knows its own ID), then every process can learn the IDs of all other processes in $n$ rounds. If input values are attached to these IDs, then the network can compute any function of such input values within $n$ rounds.
@@
Anonymous networks. Networks without IDs, also called anonymous systems, are a considerably more challenging scenario. In this model, all processes have identical initial states, and may only differ by their inputs. In the last thirty years, a large body of works have investigated the computational power of anonymous static networks, giving characterizations of what can be computed in various settings [8, 9, 13, 14, 15, 28, 52, 56].
@@
Networks with leaders. In order to deterministically solve several non-trivial problems in anonymous systems, it is necessary to have some form of initial “asymmetry” [1, 9, 42, 56]. The most common assumption is the existence of a single distinguished process called leader [2, 3, 5, 6, 22, 28, 31, 33, 42, 51, 57] or, less commonly, a subset of several leaders (and knowledge of their number [^3] It is easy to see that a network with an unknown number of leaders is equivalent to a network with no leaders at all. Also, if the leaders are distinguishable from each other, then any one of them can be elected as a unique leader. Hence, the only genuinely interesting multi-leader case is the one with a known number of indistinguishable leaders.[32, 34, 35, 36]. The presence of leaders is a realistic assumption: examples include base stations in mobile networks, gateways in sensor networks, etc. For these reasons, the computational power of anonymous systems enriched with one or more leaders has been extensively studied in the traditional model of static networks [28, 51, 57], as well as in population protocols [2, 3, 5, 6, 22].
@@
Leaderless networks. In some networks, the presence of reliable leaders may not always be guaranteed. For example, in a mobile sensor network deployed by an aircraft, the leaders may be destroyed as a result of a bad landing; also, the leaders may malfunction during the system’s lifetime. This justifies the extensive existing literature on networks with no leaders [16, 44, 45, 48, 55, 58]. Notably, a large portion of works on leaderless networks have focused on the Average Consensus problem, where the goal is to compute the mean of a list of numbers distributed among the processes [7, 26, 49, 50].
Problem classes. Assume that each process is assigned an input at the beginning of the computation, i.e., at round 00. The Input Frequency function is the function that returns the percentage of processes that have each input value. Being able to compute the Input Frequency function allows a system to compute a wide class of functions called frequency-based with no loss in performance [30]. Thus, we say that the Input Frequency function is complete for this class of functions (refer to Section 2.2). The most prominent representative of this class of functions is given by the aforementioned Average Consensus problem, since the mean of a multiset of numeric input values is a frequency-based function (the percentage of processes that have each input can be used as weight to compute the mean of all inputs).
The Input Multiset function is the function that returns the number of processes that have each input value. This function is complete for a class of functions called multiset-based, which strictly includes the frequency-based ones (see Section 2.2 for definitions). A well-studied example of a multiset-based function that is not frequency-based is given by the Counting problem, which asks to determine the total number of processes in the system. The Counting problem is practically interesting in real-world scenarios such as large-scale ad-hoc sensor networks, where the individual processes may be unaware of the size of the system [47].
@@ 1.2 State of the Art
Average Consensus problem. In the Average Consensus problem, each process starts with an input value, and the goal is to compute the average of these initial values. The typical approach found in the literature is based on local averaging algorithms, where each process updates its local value at each round based on a convex combination of the values of its neighbors. This and similar techniques for Average Consensus have been studied for decades by the distributed control and distributed computing communities  [7, 16, 34, 44, 48, 49, 50, 55, 58]. Existing works can be divided into those that give convergent solutions, those that give stabilizing solutions, and those that give terminating solutions.
In convergent algorithms, the consensus is not reached in finite time, but each process’ local value asymptotically converges to the average. A state-of-the-art $\epsilon$-convergent algorithm based on Metropolis rules with a running time of $O\left(\tau n^{3}\log(1/{\epsilon})\right)$ communication rounds is given in [44]. However, this algorithm rests on the assumption that the degree of each process in the network has a known upper bound. [^4] More generally, averaging algorithms based on Metropolis rules cannot be applied to our network model, because they require all processes to know their outdegree at every round prior to sending their messages. The running time can be improved to $O\left(n^{2}\log(1/\epsilon)\right)$ communication rounds if the network is always connected (i.e., $\tau=1$) and may only change every three rounds (as opposed to every round).
In stabilizing algorithms, the consensus is reached in a finite number of rounds, but no termination criteria are specified, and therefore all processes are assumed to run indefinitely. To the best of our knowledge, all stabilizing Average Consensus algorithms for leaderless networks are either probabilistic or assume the network to be static [26, 30, 48, 58]. However, some of these algorithms stabilize in a linear number of communication rounds [48, 58].
@@
Counting problem. A long series of papers have focused the Counting problem in connected anonymous networks with a unique leader [12, 17, 18, 19, 20, 21, 31, 32, 36, 42]. [^5] It is folklore that the Counting problem cannot be solved in leaderless networks. These works have achieved better and better running times, with the first polynomial-time algorithm presented in J. ACM in 2020, having a running time of $O\left({n^{5}}\log^{2}(n)\right)$ rounds [33]. This was later improved and extended to networks with $\ell\geq 1$ leaders in [36], solving the Counting problem in $O\left({n^{4+\epsilon}}\log^{3}(n)/\ell\right)$ communication rounds (for $\epsilon>0$).
@@
Almost all of these works share the same basic approach of implementing a mass-distribution mechanism similar to the local averaging used to solve the Average Consensus problem. [^6] We remark that the mass-distribution approach used for Counting and Average Consensus requires processes to communicate numbers whose representation size grows at least linearly with the number of rounds. Thus, all cited algorithms require processes to exchange messages of polynomial size. In spite of the technical sophistication of this line of research, there is still a striking gap in terms of running time between state-of-the-art algorithms for anonymous networks and networks with unique IDs. The same gap exists with respect to static anonymous networks, where the Counting problem is known to be solvable in $O(n)$ rounds [42]. Given the current state of the art, solving non-trivial problems in large-scale dynamic networks is still impractical.
@@ 1.3 Our Contributions
Table 1: Summary of the results in this paper. The variable $n$ indicates the number of processes in the system, $\ell$ is the number of leaders, $N$ is an upper bound on $n$, $\tau$ is the dynamic disconnectivity of the network, and $d$ is (an upper bound on) its dynamic diameter. If $\ell=0$, only the frequency-based functions are computable; if $\ell\geq 1$, the multiset-based functions are also computable.
@@
Technique. Our approach departs radically from the mass-distribution and averaging techniques traditionally adopted by most previous works on anonymous dynamic networks. Instead, all of our results are based on a novel combinatorial structure called history tree, which completely represents an anonymous dynamic network and naturally models the idea that processes can be distinguished if and only if they have different “histories” (see Section 3). Thanks to the simplicity of our technique, this paper is entirely self-contained, our proofs are transparent and easy to understand, and our algorithms are elegant and straightforward to implement.
@@
An implementation can be found at https://github.com/viglietta/Dynamic-Networks. The repository includes a dynamic network simulator that can be used to test our algorithms and visualize the history trees of custom networks.
@@
Computability. We give an exact characterization of the set of functions that can be computed deterministically in anonymous dynamic networks with and without leaders, respectively. Namely, in networks with at least one leader, a function is computable if and only if it is multiset-based; [^7] Another way of stating this result is that it is sufficient to know the size of any subset of distinguished processes in order to compute all multiset-based functions. in networks with no leaders, a function is computable if and only if it is frequency-based. [^8] A similar result, limited to static leaderless networks and functions with additional constraints, was obtained in [30]. Interestingly, computability is independent of the dynamic disconnectivity $\tau$.
@@
Algorithms. Furthermore, we give efficient deterministic [^9] An advantage of allowing randomization in our model would be the possibility of choosing unique IDs with high probability. Even though this would greatly simplify most problems, it would not afford an asymptotic improvement on the running times of our algorithms; see [38]. algorithms for computing the Input Frequency function in leaderless networks (Sections 4.2 and 4.3) and the Input Multiset function in networks with leaders (Sections 4.4 and 5). Since these functions are complete for the class of frequency-based and multiset-based functions, respectively, we automatically obtain efficient algorithms for computing all functions in these classes (Section 2.2).
@@
For each of the aforementioned functions, we give two algorithms: a terminating version, where all processes are required to commit to their output and never change it, and a slightly more efficient stabilizing version, where processes are allowed to modify their outputs, provided that they eventually stabilize on the correct output. The running times of our algorithms are summarized in Table 1.
@@
The stabilizing algorithms for both functions give the correct output within $2\tau n$ communication rounds regardless of the number of leaders, and do not require any knowledge of the dynamic disconnectivity $\tau$ or the number of processes $n$. Our terminating algorithm for leaderless networks runs in $\tau(n+N)$ communication rounds with knowledge of $\tau$ and an upper bound $N\geq n$; the terminating algorithm for $\ell\geq 1$ leaders runs in $\tau(\ell^{2}+\ell+1)n$ communication rounds [^10] Note that the case where all processes are leaders is not equivalent to the case with no leaders, because processes do not have the information that $\ell=n$, and have to “discover” that there are no non-leader processes in the network. with no knowledge of $n$. The latter running time is reasonable (i.e., linear) in most applications, as $\ell$ is typically a constant or negligible compared to $n$.
@@
We remark that the local computation time and the amount of processes’ internal memory required by our terminating algorithms is only polynomial in the size of the network. Also, like in previous works on anonymous dynamic networks, processes need to send messages of polynomial size.
Negative results. Some of our algorithms assume processes to have a-priori knowledge of some parameters of the network; in Section 6 we show that none of these assumptions can be removed. We also provide lower bounds that asymptotically match our algorithms’ running times, assuming that the number of leaders $\ell$ is constant (which is a realistic assumption in most applications). Table 1 summarizes our lower bounds, as well. We point out that our lower bounds of roughly $2\tau n$ rounds are the first non-trivial lower bounds for simple undirected anonymous dynamic networks (i.e., better than $n-1$).
Multigraphs. All of our results hold more generally if networks are modeled as multigraphs, as opposed to the simple graphs traditionally encountered in nearly all of the existing literature. This is relevant in many applications: in radio communication, for instance, multiple links between processes naturally appear due to the multi-path propagation of radio waves. Furthermore, this turns out to be a remarkably powerful feature in light of Proposition 2.4, which establishes a relationship between multigraphs and $\tau$-union-connected networks. This finding single-handedly allows us to generalize our algorithms to disconnected networks at the cost of a mere factor of $\tau$ in their running times, which is worst-case optimal.
@@
  • Knowledge of the processes. Our algorithms assume that the processes have a-priori knowledge about certain properties of the network only when the absence of such knowledge would render the Average Consensus or Counting problems unsolvable.
@@
Contribution 1. In anonymous dynamic leaderless networks, any problem that can be solved deterministically (including Average Consensus) has an algorithm that terminates (thus stabilizes) in linear time.

### v8

@@ 1.1 Background
Dynamic networks. The study of theoretical and practical aspects of highly dynamic distributed systems has garnered significant attention in recent years [12, 46, 51]. These systems involve a constantly changing network of computational devices called agents (sometimes referred to as “processes” or “processors”). Pairs of agents may send each other messages only through links that appear or disappear unpredictably. This dynamicity is typical of modern real-world systems and is the result of technological innovations, such as the spread of mobile devices, software-defined networks, wirelesses sensor networks, wearable devices, smartphones, etc.
@@
Connected networks. There are several models of dynamicity [12]; a popular choice is the 1-interval-connected network model [44, 55]. Here, a fixed set of $n$ agents communicate through links forming a time-varying graph, i.e., a graph whose edge set changes at discrete time units called rounds (thus, the system is synchronous); such a graph changes unpredictably, but is assumed to be connected at all times.
@@
Disconnected networks. The 1-interval-connected network model may not be a suitable choice for many real systems, due to the very nature of dynamic entities (think of P2P networks of smart devices moving unpredictably) or due to transient communication failures, which may compromise the network’s connectivity. A more realistic assumption is that the union of all the network’s links across any $\tau$ consecutive rounds induces a connected graph on the agents [42, 52, 57]. We say that such a network is $\tau$-union-connected, and the smallest such parameter $\tau\geq 1$ is called dynamic disconnectivity. Another widely used parameter for dynamic networks is the dynamic diameter $d$. The relationship between $\tau$ and $d$ is discussed at the end of Section 2.3. In particular, it is worth noting that every occurrence of the parameter $\tau$ in the running times of our algorithms can be safely replaced with $d$. Observe that 1-interval-connected networks can equivalently be characterized as 1-union-connected networks. We remark that non-trivial (terminating) computation requires some conditions on temporal connectivity to be met, as well as some a-priori knowledge by agents (refer to Propositions 2.3 and 2.4).
@@
Networks with unique IDs. A large number of research papers have considered dynamic systems where each agent has a distinct identity (unique IDs[29, 34, 44]. In this setting, there are efficient algorithms for consensus [45], broadcast [11], counting [44, 55], and many other fundamental problems [43, 51].
@@
It should be noted that networks with unique IDs allow for very simple algorithms for a wide variety of problems. Indeed, in a 1-interval-connected network of $n$ agents, if each agent broadcasts the set of its “known IDs” at every round (initially, an agent only knows its own ID), then every agent can learn the IDs of all other agents in $n$ rounds. If input values are attached to these IDs, then the network can compute any function of such input values within $n$ rounds.
For this reason, most research on networks with unique IDs has focused on the “congested” model, which limits the size of every message to $O(\log n)$ bits [59]. In this model, the all-to-all token dissemination problem (where each agent begins with a unique token, and the objective is for all agents to collect every token) can be solved in $O(n^{2})$ rounds using a token forwarding strategy, where tokens are merely stored, copied, and sent without any alteration [44]. Notably, any token-forwarding solution to the all-to-all token dissemination problem requires $\Omega(n^{2}/\log n)$ rounds [29].
@@
Anonymous networks. Networks without IDs, also called anonymous systems, pose an additional challenge compared to networks with unique IDs. In this model, all agents have identical initial states, and may only differ by their inputs. In the last thirty years, a large body of works have investigated the computational power of anonymous static networks, giving characterizations of what can be computed in various settings [9, 10, 14, 15, 16, 32, 61, 66].
@@
We remark that, while in congested networks with unique IDs the difficulty is to cope with the severely limited message sizes, in (non-congested) anonymous networks the difficulty is to overcome the symmetry introduced by the anonymity of agents. Hence, the tools used in these two settings are radically different.
Randomization. One may wonder how randomization could help in breaking symmetry in anonymous networks. A simple strategy would be to sample unique IDs from a source of randomness, and then leverage such unique IDs. However, this approach would require an estimate on the network size to succeed with high probability, and is therefore unsuitable for safety-critical systems (where failures are not allowed, no matter how unlikely) or systems where reliable sources of randomness are not available, or when nothing is known about the network size. Furthermore, assigning unique IDs to agents compromises their anonymity, thereby forfeiting the security and privacy benefits of having an anonymous network.
Thus, in this paper we do not consider randomized algorithms, but only deterministic ones. The only work we are aware of that proposes a randomized algorithm for determining the size of the network $n$ in the model considered in this paper is [44]. However, this algorithm only provides approximate solutions and relies on prior knowledge of an upper bound on $n$. Additionally, its correctness is guaranteed only with high probability.
The problem of estimating the network size using randomized protocols has also been studied in the peer-to-peer community [48, 36, 49]. We note that in these models, exact counting is not possible because agents continuously join and leave the network.
Networks with leaders. In order to deterministically solve several non-trivial problems in anonymous systems, it is necessary to have some form of initial “asymmetry” [1, 10, 50, 66]. The most common assumption is the existence of a single distinguished agent called leader [2, 3, 6, 7, 25, 32, 37, 39, 50, 60, 67] or, less commonly, a subset of several leaders, and knowledge of their number [38, 40, 41, 42].
Note that a network with an unknown number of leaders is equivalent to a network with no leaders at all. Also, if the leaders are distinguishable from each other, then any one of them can be elected as a unique leader. Hence, the only genuinely interesting multi-leader case is the one with a known number of indistinguishable leaders.
The presence of leaders is a realistic assumption: examples include base stations in mobile networks, gateways in sensor networks, etc. For these reasons, the computational power of anonymous systems enriched with one or more leaders has been extensively studied in the traditional model of static networks [32, 60, 67], as well as in population protocols [2, 3, 6, 7, 25].
@@
Leaderless networks. In some networks, the presence of reliable leaders may not always be guaranteed. For example, in a mobile sensor network deployed by an aircraft, the leaders may be destroyed as a result of a bad landing; also, the leaders may malfunction during the system’s lifetime. This justifies the extensive existing literature on networks with no leaders [18, 19, 52, 53, 56, 64, 68]. Notably, a large portion of works on leaderless networks have focused on the Average Consensus problem, where the goal is to compute the mean of a list of numbers distributed among the agents [8, 17, 18, 30, 57, 58].
Problem classes. Assume that each agent is assigned an input at the beginning of the computation, i.e., at round 00. The Input Frequency function is the function that returns the percentage of agents that have each input value. Being able to compute the Input Frequency function allows a system to compute a wide class of functions called frequency-based with no loss in performance [35]. Thus, we say that the Input Frequency function is complete for this class of functions (refer to Section 2.2). The most prominent representative of this class of functions is given by the aforementioned Average Consensus problem, since the mean of a multiset of numeric input values is a frequency-based function (the percentage of agents that have each input can be used as weight to compute the mean of all inputs).
The Input Multiset function is the function that returns the number of agents that have each input value. This function is complete for a class of functions called multiset-based, which strictly includes the frequency-based ones (see Section 2.2 for definitions). A well-studied example of a multiset-based function that is not frequency-based is given by the Counting problem, which asks to determine the total number of agents in the system. The Counting problem is practically interesting in real-world scenarios such as large-scale ad-hoc sensor networks, where the individual agents may be unaware of the size of the system [55].
@@ 1.2 State of the Art
Average Consensus problem. In the Average Consensus problem, each agent starts with an input value, and the goal is to compute the average of these initial values. The typical approach found in the literature is based on local averaging algorithms, where each agent updates its local value at each round based on a convex combination of the values of its neighbors. This and similar techniques for Average Consensus have been studied for decades by the distributed control and distributed computing communities  [8, 17, 18, 19, 40, 52, 56, 57, 58, 64, 68]. Existing works can be divided into those that give convergent solutions, those that give stabilizing solutions, and those that give terminating solutions.
In convergent algorithms, the consensus is not reached in finite time, but each agent’s local value asymptotically converges to the average. A state-of-the-art $\epsilon$-convergent algorithm based on Metropolis rules with a running time of $O\left(\tau n^{3}\log(1/{\epsilon})\right)$ communication rounds is given in [52]. However, this algorithm rests on the assumption that the degree of each agent in the network has a known upper bound; more generally, averaging algorithms based on Metropolis rules cannot be applied to our network model, because they require all agents to know their outdegree at every round prior to sending their messages. The running time of the algorithm in [52] can be improved to $O\left(n^{2}\log(1/\epsilon)\right)$ communication rounds if the network is always connected (i.e., $\tau=1$) and may only change every three rounds (as opposed to every round).
In stabilizing algorithms, the consensus is reached in a finite number of rounds, but no termination criteria are specified, and therefore all agents are assumed to run indefinitely. To the best of our knowledge, all stabilizing Average Consensus algorithms for leaderless networks are either probabilistic or assume the network to be static [30, 35, 56, 68]. However, some of these algorithms stabilize in a linear number of communication rounds [56, 68].
@@
Counting problem. A long series of papers have focused the Counting problem in connected anonymous networks with a unique leader [13, 20, 21, 22, 23, 24, 37, 38, 42, 50]. (We remark that the Counting problem cannot be solved in leaderless networks; this fact was implied by Angluin in [1] and is now folklore.) These works have achieved better and better running times, with the first polynomial-time algorithm presented in J. ACM in 2020, having a running time of $O\left({n^{5}}\log^{2}(n)\right)$ rounds [39]. This was later improved and extended to networks with $\ell\geq 1$ leaders in [42], solving the Counting problem in $O\left({n^{4+\epsilon}}\log^{3}(n)/\ell\right)$ communication rounds (for $\epsilon>0$).
@@
Almost all of these works share the same basic approach of implementing a mass-distribution mechanism similar to the local averaging used to solve the Average Consensus problem. We should point out that such a mass-distribution approach requires agents to communicate numbers whose representation size grows at least linearly with the number of rounds; thus, all cited algorithms require agents to exchange messages of polynomial size. In spite of the technical sophistication of this line of research, there is still a striking gap in terms of running time between state-of-the-art algorithms for anonymous networks and networks with unique IDs. The same gap exists with respect to static anonymous networks, where the Counting problem is known to be solvable in $O(n)$ rounds [50]. Given the current state of the art, solving non-trivial problems in large-scale dynamic networks is still impractical.
@@ 1.3 Our Contributions
Problem class Leaders Terminating Assumptions Running time Lower bound
Frequency-based (e.g., Average Consensus) $\ell=0$ $\tau(2n-2)$ $\tau(2n-6)$
$\tau$ and $N\geq n$ known $\tau(n+N-2)+1$ $\tau(2n-4)$
$d$ known $\tau(n-1)+d+1\leq dn+1$
[…]
$\ell$ and $\tau$ known $\tau((\ell^{2}+\ell+1)(n-1)+1)$ $\tau(2n-\ell-3)$
Table 1: Summary of the results in this paper. The variable $n$ indicates the number of agents in the system, $\ell$ is the number of leaders, $N$ is an upper bound on $n$, $\tau$ is the dynamic disconnectivity of the network (or an upper bound thereof), and $d$ is its dynamic diameter (or an upper bound thereof). If $\ell=0$, only the frequency-based functions are computable; if $\ell\geq 1$, the multiset-based functions are also computable.
@@
Technique. Our approach departs radically from the mass-distribution and averaging techniques traditionally adopted by most previous works on anonymous dynamic networks. Instead, all of our results are based on a novel combinatorial structure called history tree, which completely represents an anonymous dynamic network and naturally models the idea that agents can be distinguished if and only if they have different “histories” (see Section 3). Thanks to the simplicity of our technique, this paper is entirely self-contained, our proofs are transparent and easy to understand, and our algorithms are elegant and straightforward to implement.
In Section 3.3 we will compare history trees with other structures commonly used in the analysis of static networks. A more in-depth comparison is also found in [65]. In particular, it should be noted that from the history tree of a static network it is possible to construct the views of Yamashita–Kameda, and vice versa. Hence, a history tree encodes all the information that the agents can extract from the network.
@@
An implementation of most of the concepts discussed in this paper can be found at https://github.com/viglietta/Dynamic-Networks. The repository includes a dynamic network simulator that can be used to test our algorithms and visualize the history trees of custom networks. A browser version of the same program is also available at https://giovanniviglietta.com/projects/anonymity.
@@
Computability. We give an exact characterization of the set of functions that can be computed deterministically in anonymous dynamic networks with and without leaders, respectively. Namely, in networks with at least one leader, a function is computable if and only if it is multiset-based (another way of stating this result is that it is sufficient to know the size of any subset of distinguished agents in order to compute all multiset-based functions); on the other hand, in networks with no leaders, a function is computable if and only if it is frequency-based (a similar result, limited to static leaderless networks and functions with additional constraints, was obtained in [35]).
Observe that computability is independent of the dynamic disconnectivity $\tau$. Notably, this is a consequence of the fact that our algorithms work for multigraphs, as opposed to simple graphs, as discussed later in this section.
@@
Algorithms. Furthermore, we give efficient deterministic algorithms for computing the Input Frequency function in leaderless networks (Sections 4.2 and 4.3) and the Input Multiset function in networks with leaders (Sections 4.4 and 5). Since these functions are complete for the class of frequency-based and multiset-based functions, respectively, we automatically obtain efficient algorithms for computing all functions in these classes (Section 2.2).
@@
For each of the aforementioned functions, we give two algorithms: a terminating version, where all agents are required to commit to their output and never change it, and a slightly more efficient stabilizing version, where agents are allowed to modify their outputs, provided that they eventually stabilize on the correct output. The running times of our algorithms are summarized in Table 1.
@@
The stabilizing algorithms for both functions give the correct output within $2\tau n$ communication rounds regardless of the number of leaders, and do not require any knowledge of the dynamic disconnectivity $\tau$ or the number of agents $n$. Our terminating algorithm for leaderless networks runs in $\tau(n+N)$ communication rounds with knowledge of $\tau$ and an upper bound $N\geq n$; the terminating algorithm for $\ell\geq 1$ leaders runs in $\tau(\ell^{2}+\ell+1)n$ communication rounds with no knowledge of $n$. The latter running time is reasonable (i.e., linear) in most applications, as $\ell$ is typically a constant or negligible compared to $n$. Note that the case where all agents are leaders is not equivalent to the case with no leaders: in the former case, agents do not have the information that $\ell=n$, and have to “discover” that there are no non-leader agents in the network.
@@
We remark that the local computation time and the amount of internal memory required by our terminating algorithms is only polynomial in the size of the network. Also, like in previous works on anonymous dynamic networks, agents need to send messages of polynomial size.
Negative results. Some of our algorithms assume agents to have a-priori knowledge of some parameters of the network. In Section 6 we show that none of these assumptions can be removed, in the sense that removing any one of them would make the corresponding problem unsolvable by any algorithm. Note that we are not implying that this specific a-priori knowledge is strictly necessary in an absolute sense: it remains possible that alternative information, different from what we assume, could also suffice.
We also provide lower bounds that asymptotically match our algorithms’ running times, assuming that the number of leaders $\ell$ is a constant (which is a realistic assumption in most applications). Table 1 summarizes our lower bounds, as well. We point out that our lower bounds of roughly $2\tau n$ rounds are the first non-trivial lower bounds for simple undirected anonymous dynamic networks (i.e., better than $n-1$).
Multigraphs. All of our results hold more generally if networks are modeled as multigraphs, as opposed to the simple graphs traditionally encountered in nearly all of the existing literature. This is relevant in many applications: in radio communication, for instance, multiple links between agents naturally appear due to the multi-path propagation of radio waves.
Furthermore, this turns out to be a remarkably powerful feature in light of Proposition 2.5, which establishes a relationship between multigraphs and $\tau$-union-connected networks. This finding single-handedly allows us to generalize our algorithms to disconnected networks at the cost of a mere factor of $\tau$ in their running times, which is worst-case optimal.
It is worth remarking that, while our algorithms apply to multigraphs, all the impossibility results and counterexamples in Section 6 hold for simple graphs. Therefore, all of our results can be stated equivalently for both multigraphs and simple graphs.
@@
  • Knowledge of the agents. Our algorithms assume that the agents have a-priori knowledge about certain properties of the network only when the absence of such knowledge would render the Average Consensus or Counting problems unsolvable.
@@
Contribution 1. In anonymous dynamic leaderless networks, any problem that can be solved deterministically (including Average Consensus) has an algorithm that stabilizes in linear time. If an upper bound on the number of agents is known, the algorithm also terminates in linear time.